# Pruning via Merging: Compressing LLMs via Manifold Alignment Based Layer Merging

### v1

@@ 2.2 Similarity-based Layer Merging
1:Input: LLM $\mathcal{M}$ with Layers ${L_{1},L_{2},\dots,L_{N}}$, Layer Parameters $\Theta={\theta_{1},\theta_{2},\dots,\theta_{N}}$, Dataset $\omega$
2:Output: Compressed Model $\mathcal{M}^{*}$ with Aligned Knowledge
3:$\mathcal{H}\leftarrow\text{ExtractActivations}(\mathcal{M},\omega)$triangleright Extract activations for each layer on dataset $\omega$
4:$D\leftarrow\text{ComputePairwiseDistances}(\mathcal{H})$triangleright Compute pairwise Euclidean distance matrix of activations
5:$\mathbf{E}\leftarrow\text{DiffusionKernel}(D,\sigma_{K})$triangleright Apply diffusion kernel for manifold learning
6:$\mathcal{S}\leftarrow\text{ComputeNPIB}(\mathbf{E})$triangleright Compute NPIB similarity matrix
7:$\Omega\leftarrow\text{SortLayersBySimilarity}(\mathcal{S})$triangleright Sort layers by similarity for merging
8:while $|\Omega|>1$ do
9:     $(L_{i},L_{j})\leftarrow\text{SelectTopLayerPair}(\Omega)$triangleright Select top-ranked layer pair based on similarity
10:     $\lambda_{m}\leftarrow\text{ComputeMergingRatio}(\mathcal{S},L_{i},L_{j})$triangleright Compute adaptive merging ratio
11:     $\theta_{m}\leftarrow\lambda_{m}\cdot\theta_{i}+(1-\lambda_{m})\cdot\theta_{j}$triangleright Merge layer parameters
12:     $L_{m}\leftarrow\text{MergeLayer}(L_{i},L_{j},\theta_{m})$triangleright Create merged layer using the merged parameters
13:     $\mathcal{M}\leftarrow\text{ReplaceLayer}(\mathcal{M},L_{i},L_{j},L_{m})$triangleright Update model with the merged layer
14:     $\Omega\leftarrow\text{UpdateLayerList}(\Omega,L_{i},L_{j},L_{m})$triangleright Update layer list with the new merged layer
@@
16:return $\mathcal{M}$
Building upon the manifold learning representations, MKA employs a similarity-based layer merging approach to identify and fuse layers with highly aligned knowledge. By quantifying the similarity between layers using the Normalized Pairwise Information Bottleneck (NPIB) Tishby et al. (2000) metric, we can determine which layers are most suitable for merging. This process allows us to reduce model size, improve inference speed, and decrease GPU memory consumption.
The layer merging process involves several key steps. First, we construct a similarity matrix using the NPIB metric to compare the knowledge patterns across layers. Next, we introduce an adaptive weight allocation function to determine the optimal merging ratio for each pair of layers, ensuring that the merged layer retains the most critical features. Finally, we fuse the parameters of the selected layers using the weighted sum and update the model architecture accordingly.
Constructing the Similarity Matrix: To identify layers suitable for merging, we first construct a similarity matrix $S$ using the Normalized Pairwise Information Bottleneck (NPIB) metric. NPIB quantifies the shared information between layers while normalizing their individual entropies, providing an ideal measure for comparing knowledge patterns across layers:
$S_{ij}$ $=\text{NPIB}(\mathbf{E}_{i},\mathbf{E}_{j})$ (3)
$=\frac{\sum\limits_{x\in\mathbf{E}_{i}}\sum\limits_{y\in\mathbf{E}_{j}}p(x,y)\log\frac{p(x,y)}{p(x)p(y)}}{\sqrt{\sum\limits_{x\in\mathbf{E}_{i}}p(x)\log p(x)\cdot\sum\limits_{y\in\mathbf{E}_{j}}p(y)\log p(y)}}$
where $p(x,y)$ denotes the joint probability distribution of $\mathbf{E}_{i}$ and $\mathbf{E}_{j}$, and $p(x)$ and $p(y)$ represent the marginal probability distributions of $\mathbf{E}_{i}$ and $\mathbf{E}_{j}$, respectively. This similarity matrix helps us determine which layers have the most aligned knowledge representations.
Calculate Weight ratio: To determine the merging ratio $\lambda_{m}$ for each pair of layers, we introduce the adaptive weight allocation function $\Psi$. This function dynamically adjusts the merging ratio based on the similarity differences between layers, ensuring that the merged layer retains the most critical features from each original layer:
$\lambda_{m}=\Psi(\bar{s}i,\bar{s}j)=\frac{e^{\mathcal{S}{ij}}}{\sum{k\in\Omega}e^{\mathcal{S}_{k}}}$ (4)
The adaptive weight allocation function $\Psi$ adjusts the merging weights based on the similarity difference between layers. When the similarity difference between two layers is large, $\Psi$ assigns a higher weight to the layer with higher similarity, reducing the weight of the layer with lower similarity. This mechanism ensures that the merged layer better preserves the knowledge from the more similar layer.
Merging Layer Parameters: Once the merging ratio $\lambda_{m}$ is determined, we fuse the parameters $\mathbf{\theta}_{i}$ and $\mathbf{\theta}_{j}$ of the selected layers using a weighted sum:
$\widetilde{\mathbf{\theta}}_{m}=\lambda_{m}\mathbf{\theta}_{i}+(1-\lambda_{m})\mathbf{\theta}_{j}$ (5)
The merged layer $L_{m}$ is obtained through the function $\text{FuseLayer}(L_{i},L_{j},\widetilde{\mathbf{\theta}}_{m})$, which constructs a new layer based on the fused parameters $\widetilde{\mathbf{\theta}}_{m}$. This new layer integrates the aligned knowledge from the original layers, preserving essential information while reducing redundancy.
Finally, we update the model $\mathcal{M}$ by replacing the original layers $L_{i}$ and $L_{j}$ with the newly merged layer $L_{m}$, utilizing the function $\text{ReplaceLayer}(\mathcal{M},L_{i},L_{j},L_{m})$. This step ensures that the model’s architecture is updated to reflect the compression process, maintaining performance while significantly reducing model size.
Figure 2: Comparison of Accuracy (ACC) during merging and pruning on the MMLU dataset. MKA achieves higher compression ratios (approximately 43.5% for Llama3-8B, 45% for Llama3-70B, 40% for Mistral-7B, 31.25% for Llama2-7B, and 57.5% for Llama2-13B) while preserving 90% performance. Please see the appendix A for details.
@@ 3.1.1 Datasets
We conduct evaluations using the MKA methods across various benchmark datasets, each specifically designed to test various facets of language comprehension and generation. In detail, MMLU Hendrycks et al. (2020) evaluates broad language understanding across a wide range of domains. PIQA Bisk et al. (2020) is designed to test models on commonsense reasoning in the physical world, aiming to assess NLP models’ grasp of everyday physical interactions. HellaSwag Zellers et al. (2019) is a challenge dataset for commonsense natural language inference, consisting of event descriptions with multiple possible continuations, where the task is to select the most plausible one. RACE-H Lai et al. (2017) is a large-scale reading comprehension dataset collected from English exams for Chinese high school students, featuring a high proportion of questions that require reasoning. BoolQ Clark et al. (2019) is a reading comprehension dataset focusing on naturally occurring yes/no questions that often query for complex, non-factoid information and require difficult entailment-like inference to answer correctly.
@@ 3.1.2 LLMs
In our experiments, we employ the Llama-2 Touvron et al. (2023), Llama-3, and Mistral-7B Jiang et al. (2023) models, each distinct in their capabilities and configurations: Llama-2: Encompassing models from 7 billion to 70 billion parameters, exhibits superior performance and safety on diverse benchmarks. Llama-3: Featuring models with 8 billion and 70 billion parameters, Llama3 offers state-of-the-art performance and advanced reasoning capabilities. Mistral-7B: a 7-billion-parameter model that surpasses Llama-2 and Llama-1 in performance and efficiency, leveraging grouped-query and sliding window attention mechanisms for optimal inference across lengthy sequences.
@@ 3.1.3 Baselines
In this study, we assess the effectiveness of our proposed method, MKA, through two distinct comparative analyses. Firstly, we evaluate MKA directly against several well-established pruning techniques to gauge its standalone efficacy in reducing model size while maintaining performance. Secondly, we extend the comparison to include scenarios where both the traditional pruning methods and MKA are further enhanced through quantization. The baseline methods included in our analysis are: SparseGPT Frantar and Alistarh (2023): An efficient one-shot pruning method that can induce high sparsity levels in large language models with billions of parameters while preserving accuracy, by reducing the pruning problem to a set of large-scale sparse regression instances solved by a novel approximate solver. ShortGPT Men et al. (2024): A pruning method that removes redundant layers from large language models based on a Block Influence metric, which assesses the significance of each layer. Reverse Pruning: A heuristic approach where the importance of layers is considered inversely proportional to their order in the model, prioritizing the retention of earlier layers. SmoothQuant Xiao et al. (2023): SmoothQuant is a training-free post-training quantization solution that enables efficient 8-bit weight and activation quantization for large language models, offering up to 1.56× speedup and 2× memory reduction with minimal accuracy loss. GPTQ Frantar et al. (2022): A one-shot weight quantization method that uses approximate second-order information to maintain high accuracy even with severe weight reduction. AWQ Lin et al. (2023): A novel quantization approach that protects salient weights by adjusting per-channel scaling based on activation observations rather than weight Magnitudes.
3.2 In what ways does MKA surpass conventional pruning techniques?
@@
We compare the performance of MKA with baseline compression methods on the MMLU dataset using the Llama3-8B, Llama3-70B, Mistral-7B, Llama2-7B, and Llama2-13B models. The evaluation metric is Accuracy (ACC) during merging and pruning. The results are presented in Figure 2.
@@
We compare the performance of MKA with baseline compression methods on the MMLU dataset using the Llama3-8B, Llama3-70B, Mistral-7B, Llama2-7B, and Llama2-13B models. The evaluation metrics include Accuracy (ACC) during merging and pruning. The results are presented in Figure 2. We can observe that, across all models, our method improves the compression ratio while maintaining performance. Specifically, the compression ratio [^2] Note that, the compression ratio is calculated as: $\left(L_{\text{total}}-\left(\frac{L_{\text{retained}}}{Q}\right)\right)/L_{\text{total}}$, where $L_{total}$ is the total number of layers before compression, $L_{retained}$ is the number of retained layers, and $Q$ is the quantization factor. for Llama3-8B reach 43.5%, for Mistral-7B it reaches 40%, and for Llama2-13B it reaches an impressive 57.5%. Additionally, we observe several phenomena: both methods experience a collapse in model performance, but the model merging method can delay the layer collapse to some extent and stabilize the model’s performance very well. Since our strategy is based on Reverse Prune, the scores for the Llama3-8B, Llama2-7B, and Llama2-13B models are very close to the Reverse Prune. Our hypothesis is that the pruning or merging of these models is similar, but model merging can adjust the merging ratio to surpass the effect of pruning. Moreover, for the Llama3-70B and Mistral-7B models, we noticed that the results do not closely match the Reverse Prune.
@@
Llama2-7B Vanilla Model 32(0.00%) 46.67
ShortGPT+Smooth 16(87.50%) 25.67
ShortGPT+GPTQ 16(87.50%) 25.82
ShortGPT+AWQ 16(87.50%) 26.01
MKA (Ours) + Smooth 16(87.50%) 35.66 (+9.99)
MKA (Ours) + GPTQ 16(87.50%) 35.91 (+10.09)
MKA (Ours) + AWQ 16(87.50%) 36.23 (+10.22)
Llama2-13B Vanilla Model 40 (0.00%) 55.62
ShortGPT+Smooth 20 (87.50%) 25.89
ShortGPT+GPTQ 20 (87.50%) 25.35
ShortGPT+AWQ 20 (87.50%) 23.83
MKA (Ours) + Smooth 20 (87.50%) 46.82 (+20.93)
MKA (Ours) + GPTQ 20 (87.50%) 45.44 (+20.09)
MKA (Ours) + AWQ 20 (87.50%) 45.86 (+22.03)
Table 1: Performance comparison of MKA and ShortGPT pruning with quantization (SmoothQuant, GPTQ, AWQ) on MMLU using Llama3-8B, Mistral-7B, Llama2-7B, and Llama2-13B. MKA outperforms ShortGPT in accuracy across all models and quantization methods at similar compression ratios with int4. The calculation of the compression ratio only considers the number of hidden layers in the model without considering the embedding layer.
@@
MKA (Ours) 63.42 65.61 48.83 55.26 63.58 64.42 65.51 45.10 45.91 62.14
Table 2: Comparison of different methods across MMLU, PIQA, HellaSwag, RACE-H, and BoolQ datasets at different compression ratios.
Figure 3: Similarity matrices for Llama-3-8B, Llama-3-70B, Mistral-7B, Llama-2-7B, and Llama-2-13B before and after MKA. Later layers show high similarity, supporting layer merging.

### v2

@@ 4.6 Manifold-Based Knowledge Alignment and Layer Merging Compression (MKA)
0:  LLM $\mathcal{M}$ with layers $\{L_{1},L_{2},\dots,L_{N}\}$ and parameters $\Theta=\{\theta_{1},\theta_{2},\dots,\theta_{N}\}$, dataset $\omega$
0:  Compressed model $\mathcal{M}^{*}$ with aligned representations
1:  Extract activations =$\mathcal{H}=\{\mathbf{H}^{l}_{i}\}$ for each layer $l$ on dataset $\omega$
2:  for each layer $l$ do
3:     Compute pairwise distances among activations $\{\mathbf{H}^{l}_{i}\}$
4:     Construct the affinity matrix $W^{(l)}$ using a Gaussian kernel with bandwidth $\sigma$
5:     Compute the diffusion map embedding $\mathbf{\Psi}^{l}$ via the eigendecomposition of $W^{(l)}$
6:  end for
7:  for each pair of layers $(l,m)$ do
8:     Estimate the covariance matrices $\Sigma_{\mathbf{\Psi}^{l}}$, $\Sigma_{\mathbf{\Psi}^{m}}$, and joint covariance $\Sigma_{\mathbf{\Psi}^{l},\mathbf{\Psi}^{m}}$
9:     Compute $I(\mathbf{\Psi}^{l};\mathbf{\Psi}^{m})$ and the normalized similarity score $S_{lm}$
10:  end for
11:  while there exists a pair $(l,m)$ with $S_{lm}\geq\tau$ do
12:     Determine the merging weight $\alpha$, for example via $\alpha=S_{lm}$ (or, alternatively, using a softmax formulation)
13:     Merge the parameters: $\widetilde{\theta}_{c}=\alpha\,\theta_{l}+(1-\alpha)\,\theta_{m}$
14:     Replace layers $l$ and $m$ with the merged layer using $\widetilde{\theta}_{c}$ and update $\mathcal{M}$ accordingly
@@
16:  return  Compressed model $\mathcal{M}^{*}$
4.7 Layer Merging via the Information Bottleneck Principle
To guide the layer merging process, we adopt an information-theoretic perspective based on the Information Bottleneck (IB) principle. The IB framework seeks to extract a compressed representation that preserves the information most relevant to a target variable.
In its standard form, given random variables $\mathbf{X}$ and $\mathbf{Y}$, the IB objective is to find a mapping $p(\mathbf{T}|\mathbf{X})$ that minimizes
$\mathcal{L}_{\mathrm{IB}}=I(\mathbf{X};\mathbf{T})-\beta\,I(\mathbf{T};\mathbf{Y}),$ (24)
where $\beta>0$ balances the trade-off between compression (minimizing $I(\mathbf{X};\mathbf{T})$) and prediction (maximizing $I(\mathbf{T};\mathbf{Y})$).
For merging layers $l$ and $m$, we set $\mathbf{X}=(\mathbf{\Psi}^{l},\mathbf{\Psi}^{m})$ and seek a compressed representation $\mathbf{\Psi}^{c}$ that captures their shared information. The corresponding IB objective is
$\mathcal{L}_{\mathrm{IB}}=I\bigl{(}(\mathbf{\Psi}^{l},\mathbf{\Psi}^{m});\mathbf{\Psi}^{c}\bigr{)}-\beta\,I(\mathbf{\Psi}^{c};\mathbf{Y}).$ (25)
To make the optimization tractable, we restrict the mapping to a deterministic linear combination:
$\mathbf{\Psi}^{c}=\alpha\,\mathbf{\Psi}^{l}+(1-\alpha)\,\mathbf{\Psi}^{m},\quad\alpha\in[0,1].$ (26)
Under the joint Gaussian assumption, the covariance of $\mathbf{\Psi}^{c}$ is given by
$\Sigma_{\mathbf{\Psi}^{c}}=\alpha^{2}\,\Sigma_{\mathbf{\Psi}^{l}}+(1-\alpha)^{2}\,\Sigma_{\mathbf{\Psi}^{m}}+2\alpha(1-\alpha)\,\Sigma_{\mathbf{\Psi}^{l},\mathbf{\Psi}^{m}},$ (27)
and if $\Sigma_{\mathbf{\Psi}^{c},\mathbf{Y}}$ denotes the cross-covariance between $\mathbf{\Psi}^{c}$ and $\mathbf{Y}$, then
$\Sigma_{\mathbf{\Psi}^{c},\mathbf{Y}}=\alpha\,\Sigma_{\mathbf{\Psi}^{l},\mathbf{Y}}+(1-\alpha)\,\Sigma_{\mathbf{\Psi}^{m},\mathbf{Y}}.$ (28)
The conditional covariance is defined as
$\Sigma_{\mathbf{\Psi}^{c}|\mathbf{Y}}=\Sigma_{\mathbf{\Psi}^{c}}-\Sigma_{\mathbf{\Psi}^{c},\mathbf{Y}}\,\Sigma_{\mathbf{Y}}^{-1}\,\Sigma_{\mathbf{Y},\mathbf{\Psi}^{c}}.$ (29)
Since $\mathbf{\Psi}^{c}$ is a deterministic function of $(\mathbf{\Psi}^{l},\mathbf{\Psi}^{m})$ (i.e., $H(\mathbf{\Psi}^{c}|(\mathbf{\Psi}^{l},\mathbf{\Psi}^{m}))=0$), the mutual information terms reduce to differential entropies. In particular, using the standard Gaussian formula for differential entropy,
$H(\mathbf{\Psi}^{c})$ $=\frac{1}{2}\ln\Bigl{(}(2\pi e)^{d_{c}}|\Sigma_{\mathbf{\Psi}^{c}}|\Bigr{)},$ (30)
$H(\mathbf{\Psi}^{c}|\mathbf{Y})$ $=\frac{1}{2}\ln\Bigl{(}(2\pi e)^{d_{c}}|\Sigma_{\mathbf{\Psi}^{c}|\mathbf{Y}}|\Bigr{)},$ (31)
the IB objective simplifies to
$\mathcal{L}_{\mathrm{IB}}=\frac{1}{2}\Bigl{[}(1-\beta)\ln|\Sigma_{\mathbf{\Psi}^{c}}|+\beta\,\ln|\Sigma_{\mathbf{\Psi}^{c}|\mathbf{Y}}|\Bigr{]}+\mathrm{const},$ (32)
where the constant term is independent of $\alpha$.
In principle, one may optimize $\mathcal{L}_{\mathrm{IB}}$ with respect to $\alpha$ by differentiating
$\frac{\partial\mathcal{L}_{\mathrm{IB}}}{\partial\alpha}=\frac{1}{2}\Biggl{[}(1-\beta)\,\mathrm{Tr}\Bigl{(}\Sigma_{\mathbf{\Psi}^{c}}^{-1}\frac{\partial\Sigma_{\mathbf{\Psi}^{c}}}{\partial\alpha}\Bigr{)}+\beta\,\mathrm{Tr}\Bigl{(}\Sigma_{\mathbf{\Psi}^{c}|\mathbf{Y}}^{-1}\frac{\partial\Sigma_{\mathbf{\Psi}^{c}|\mathbf{Y}}}{\partial\alpha}\Bigr{)}\Biggr{]}$ (33)
and setting the derivative to zero. For example, one may compute
$\frac{\partial\Sigma_{\mathbf{\Psi}^{c}}}{\partial\alpha}=2\alpha\,\Sigma_{\mathbf{\Psi}^{l}}-2(1-\alpha)\,\Sigma_{\mathbf{\Psi}^{m}}+2(1-2\alpha)\,\Sigma_{\mathbf{\Psi}^{l},\mathbf{\Psi}^{m}}.$ (34)
A similar expression can be derived for $\partial\Sigma_{\mathbf{\Psi}^{c}|\mathbf{Y}}/\partial\alpha$. In practice, however, this equation does not admit a closed-form solution. Hence, we adopt an approximate strategy by setting
$\alpha=S_{lm},$ (35)
where $S_{lm}$ is the normalized mutual information defined earlier. This heuristic assigns greater weight to the layer with higher shared information, thereby approximately minimizing the IB objective while maintaining computational efficiency.
With the weight $\alpha$ determined, the parameters of layers $l$ and $m$ are merged via
$\widetilde{\bm{\theta}}_{c}=\alpha\,\bm{\theta}_{l}+(1-\alpha)\,\bm{\theta}_{m}.$ (36)
This merged parameter set $\widetilde{\bm{\theta}}_{c}$ is used to replace the original layers, resulting in a compressed model $\mathcal{M}^{*}$.
4.8 Impact on Model Performance
Merging layers alters the parameter space of the model, which may in turn affect its performance.
Let $\delta\bm{\theta}=\widetilde{\bm{\theta}}_{c}-\bm{\theta}$ denote the shift in parameters due to merging. Suppose that the loss function $\mathcal{L}(\bm{\theta})$ is twice differentiable and locally convex in a neighborhood around $\bm{\theta}$. A second-order Taylor expansion about $\bm{\theta}$ yields
$\mathcal{L}(\widetilde{\bm{\theta}}_{c})\approx\mathcal{L}(\bm{\theta})+\nabla\mathcal{L}(\bm{\theta})^{\top}\delta\bm{\theta}+\frac{1}{2}\,\delta\bm{\theta}^{\top}\nabla^{2}\mathcal{L}(\bm{\theta})\,\delta\bm{\theta}.$ (37)
At a local minimum where $\nabla\mathcal{L}(\bm{\theta})=\mathbf{0}$, the increase in loss is bounded by
$\Delta\mathcal{L}$ $=\mathcal{L}(\widetilde{\bm{\theta}}_{c})-\mathcal{L}(\bm{\theta})$
$\leq\frac{1}{2}\,\lambda_{\max}\,\|\delta\bm{\theta}\|^{2},$ (38)
where $\lambda_{\max}$ is the largest eigenvalue of the Hessian $\nabla^{2}\mathcal{L}(\bm{\theta})$. This bound follows from the Rayleigh quotient and provides a guarantee on the loss increase due to the merging operation.
@@ 5.1.1 Datasets
We conduct evaluations using the MKA methods across various benchmark datasets, each specifically designed to test various facets of language comprehension and generation. In detail, MMLU [46] evaluates broad language understanding across a wide range of domains. PIQA [47] is designed to test models on commonsense reasoning in the physical world, aiming to assess NLP models’ grasp of everyday physical interactions. HellaSwag [48] is a challenge dataset for commonsense natural language inference, consisting of event descriptions with multiple possible continuations, where the task is to select the most plausible one. RACE-H [49] is a large-scale reading comprehension dataset collected from English exams for Chinese high school students, featuring a high proportion of questions that require reasoning. BoolQ [50] is a reading comprehension dataset focusing on naturally occurring yes/no questions that often query for complex, non-factoid information and require difficult entailment-like inference to answer correctly.
@@ 5.1.2 LLMs
In our experiments, we employ the Llama2 [3], Llama3, Llama3.2, and Mistral [51] models, each distinct in their capabilities and configurations: Llama2: Encompassing models from 7 billion to 13 billion parameters, exhibits superior performance and safety on diverse benchmarks. Llama3: Featuring models with 8 billion parameters, which offers state-of-the-art performance and advanced reasoning capabilities. Llama3.2: Featuring models with 3 billion parameters that balances performance and number of parameters. Mistral: We use the 7 billion parameter version of Mistral that surpasses Llama-2 and Llama-1 in performance and efficiency, leveraging grouped-query and sliding window attention mechanisms for optimal inference across lengthy sequences.
@@ 5.1.3 Baselines
In this study, we assess the effectiveness of our proposed method, MKA, through two distinct comparative analyses. Firstly, we evaluate MKA directly against several well-established pruning techniques to gauge its standalone efficacy in reducing model size while maintaining performance. Secondly, we extend the comparison to include scenarios where both the traditional pruning methods and MKA are further enhanced through quantization. The baseline methods included in our analysis are: PruneMe [52]: A pruning method identifies the optimal block of layers to prune by considering the similarity across layers. SLEB [53]: A pruning method designed to streamline LLMs by eliminating redundant transformer blocks. We choose the transformer block as the fundamental unit for pruning, because LLMs exhibit block-level redundancy with high similarity between the outputs of neighboring blocks. Shortened [54]: A pruning method first uses a simple metric to identify unimportant blocks and then performs a simple one-shot pruning. ShortGPT [14]: A pruning method that removes redundant layers from large language models based on a Block Influence metric, which assesses the significance of each layer. Reverse: A heuristic approach where the importance of layers is considered inversely proportional to their order in the model, prioritizing the retention of earlier layers. SmoothQuant [55]: SmoothQuant is a training-free post-training quantization solution that enables efficient 8-bit weight and activation quantization for large language models, offering up to 1.56× speedup and 2× memory reduction with minimal accuracy loss. GPTQ [56]: A one-shot weight quantization method that uses approximate second-order information to maintain high accuracy even with severe weight reduction. AWQ [57]: A novel quantization approach that protects salient weights by adjusting per-channel scaling based on activation observations rather than weight Magnitudes.
Figure 2: Performance (Accuracy) of LLMs (Llama2-7B, Llama2-13B, Llama3-8B, Llama3.2-3B, and Mistral-7B) on the MMLU dataset as the pruning ratio of various pruning methods increases.
6 Experiments
6.1 Comparison of MKA with other structured pruning methods
@@
We compare the performance of MKA with baseline compression methods on the MMLU dataset using the Llama2-7B, Llama2-13B, Llama3-8B, Llama3.2-3B, and Mistral-7B models. The evaluation metric is Accuracy (ACC) during merging and pruning. The results are presented in Figure 2.
@@
From Figure 2, we can observe that, across all models, our method improves the compression ratio while maintaining performance. Specifically, the compression ratio [^3] Note that, the compression ratio is calculated as: $\left(L_{\text{total}}-\left(\frac{L_{\text{retained}}}{Q}\right)\right)/L_{\text{total}}$, where $L_{total}$ is the total number of layers before compression, $L_{retained}$ is the number of retained layers, and $Q$ is the quantization factor. for Llama2-7B reach 31%, for Llama3-8B reach 44%, for Llama3.2-3B reach 43%, for Mistral-7B it reaches 40%, and for Llama2-13B it reaches an impressive 58%. Additionally, we observe several phenomena: both methods experience a collapse in model performance, but the model merging method can delay the layer collapse to some extent and stabilize the model’s performance very well. Since our strategy is based on Reverse Prune, the scores for the Llama2-7B, Llama2-13B, Llama3-8B, and Llama3.2-3B models are very close to the Reverse Prune. Our hypothesis is that the pruning or merging of these models is similar, but model merging can adjust the merging ratio to surpass the effect of pruning. Moreover, for the Mistral-7B models, we noticed that the results do not closely match the Reverse Prune.
TABLE I: Performance comparison of MKA and ShortGPT pruning with quantization (SmoothQuant, GPTQ, AWQ) on MMLU using Llama2-7B, Llama3-8B, and Mistral-7B. MKA outperforms ShortGPT in accuracy across all models and quantization methods at similar compression ratios with int4. The calculation of the compression ratio only considers the number of hidden layers in the model without considering the embedding layer.
@@
Llama2-7B Vanilla Model 32(0.00%) 46.67
ShortGPT+Smooth 16(87.50%) 25.67
ShortGPT+GPTQ 16(87.50%) 25.82
ShortGPT+AWQ 16(87.50%) 26.01
MKA (Ours) + Smooth 16(87.50%) 35.66 (+9.99)
MKA (Ours) + GPTQ 16(87.50%) 35.91 (+10.09)
MKA (Ours) + AWQ 16(87.50%) 36.23 (+10.22)
@@
TABLE II: Comparison of different methods across MMLU, PIQA, HellaSwag, RACE-H, and BoolQ datasets at different compression ratios.
@@
MKA (Ours) 63.42(+24.16) 65.61(+6.25) 48.83(+14.67) 55.26(+33.20) 63.58(+1.81) 64.42(+31.31) 65.51(+6.48) 45.10(+11.35) 45.91(+22.77) 62.14(+0.51)
Figure 3: Similarity matrices for Llama2-7B, Llama2-13B, Llama-3-8B, Llama3.2-3B, and Mistral-7B before MKA. Later layers show high similarity, supporting layer merging.